b) Solve (ex+e)y=x-e using the substitution v= e (this substitution will produce a homogeneous equation).

with clear steps please.
Answer #1

given (e. are 24 y = x-edo By substituting ved in Now ③ in (v.x+v2) I, dv = x-v v de → v(x+u). + d iv x -v Xtv dva 1-v/X tva→ (2+1) dz d a (1-22-2) Integrating on both sides Now Here put 1-2z-zhat dt= -2dZ-2 Zdz. at=-21 Z+Daz - 1 Int = lnk+ Ine wher> z= -2 + J4_4[2652-1) → 7 -2+2 Si-bc)2+1 > z= + √2-452 - we know that z = = --603-3 98. -11212-2012 > y = ln (-2 ta Ja_002)